# Rethinking the Sampling Criteria in Reinforcement Learning for LLM Reasoning: A Competence-Difficulty Alignment Perspective

### v2

@@ 7 Conclusion
Appendix A Training Details
@@
Table 2 presents the key configuration used for our Qwen2.5-7B experiment. Training was conducted on a single node with 8 A100 GPUs, and the model was trained for 110 steps using the veRL library [31]. For Qwen2.5-14B experiment, training was conducted on four nodes, each equipped with 8 A100 GPUs. The model was trained for 200 steps with a batch size of 256. Other configurations were kept consistent with those used in the Qwen2.5-7B experiment.
@@
Appendix B Evaluation Benchmarks
@@
  • MATH500 [12]: A curated subset of the larger MATH dataset, consisting of 500 problems that span various mathematical fields such as algebra, geometry, and number theory.
@@
  • MinervaMath [15]: A benchmark used to evaluate the performance of large language models on detailed, multi-step quantitative reasoning problems.
@@
  • OlympiadBench [16]: A benchmark specifically designed for assessing models on Olympiad-level mathematics, including problems from national and international math competitions.
@@
  • GSM8K [17]: A dataset of grade school math word problems, commonly used to test arithmetic and step-by-step reasoning in models.
Appendix C Convergence Analysis
Recall the following coupled fixed-point system:
$\begin{cases}D^{*}(x)=\sigma(C^{*}-D^{*}(x))-S^{*}(x),\quad x\in\mathbb{X}\\ C^{*}=-\mathbb{E}_{x}[D^{*}(x)]\end{cases}$ (13)
The sigmoid function $\sigma(z)$ is Lipschitz continuous[36] with constant $L_{\sigma}=\frac{1}{4}$. For any $D(x),D^{\prime}(x)$ and $C,C^{\prime}$, we have:
Then,
Taking the maximum over all $x$, we obtain
For the update of $C$, we have
Therefore, the joint mapping satisfies.
That is, the contraction constant is $\frac{1}{2}<1$.
Since the joint mapping is a contraction in the $(N+1)$-dimensional space[37], by the Banach fixed-point theorem, the sequence $(\mathbf{D}_{n},C_{n})$ converges to a unique fixed point $(\mathbf{D}^{*},C^{*})$, regardless of the initial values.
Appendix D Full Experimental Results
LABEL:fig:7bfullresults displays the full results for the Qwen2.5-7B model, while LABEL:fig:14bfullresults displays the corresponding outcomes for the Qwen2.5-14B model.
Limitations
(a) Easy-biased dataset
(b) Hard-biased dataset
Figure 19: The impact of dataset difficulty distribution.
In our main experiments, the difficulty of the problems in MATH is distributed around the model’s capabilities. However, the difficulty distribution of the dataset may be biased in practical scenes. To explore the impact of the overall difficulty of the dataset on the training of CDAS, we sample an extremely easy subset containing 3000 problems with the highest pass rate and an extremely difficult subset containing 3000 problems with the lowest pass rate based on the results of Qwen2.5-7B. We compare CDAS with Prioritized Sampling and Random Sampling, training them for 30 steps.
We observe an interesting phenomenon that when the overall difficulty of the dataset is extremely biased, there is no obvious difference between different sampling strategies (see fig. 19). We conjecture that one of the prerequisite for sampling strategies to be effective in the RL training is that the difficulty distribution of the dataset does not deviate too much from the model’s capabilities.

### v3

@@ Conclusion
Appendix A Related Work
RL for LLMs reasoning. Reinforcement learning (RL) has been widely adopted to enhance the reasoning abilities of LLMs, especially in mathematics and programming tasks (He et al. 2024; Trinh et al. 2024; Jain et al. 2025; Penedo et al. 2025). Actor-critic-based methods, such as Proximal Policy Optimization (PPO) (Schulman et al. 2017), utilize a value model to estimate the value function, guiding the policy updates. On the other hand, REINFORCE-based methods rely on policy gradients without a value model. Group Relative Policy Optimization (GRPO) (Shao et al. 2024) normalizes rewards within a group of generated outputs, eliminating the need for a separate value model. REINFORCE++ (Hu 2025) enhances the classical REINFORCE algorithm (Williams 1992) by incorporating optimization techniques from PPO. Dynamic Sampling Policy Optimization (DAPO) (Yu et al. 2025) introduces several optimizations to enhance training efficiency and stability in long-CoT reasoning tasks.
Sampling Strategies for RL training. Effective sampling strategies are crucial for efficient RL training with LLMs. Coarse-grained curriculum learning (Team et al. 2025; Xie et al. 2025) gradually increases problem difficulty based on predefined labels. LIMR (Li et al. 2025) introduces Learning Impact Measurement (LIM) to select problems that align with the model’s learning trajectory. Prioritized Sampling (Team et al. 2025) tracks the pass rate for each problem and samples problems in proportion to their failure rates. This approach directs the model’s focus toward more challenging problems. Dynamic Sampling (Yu et al. 2025) continues to sample problems within a batch until their pass rates are neither 0 nor 1.
Appendix B Convergence Analysis
Recall the following coupled fixed-point system:
$\begin{cases}D^{*}(x)=\sigma(C^{*}-D^{*}(x))-S^{*}(x),\quad x\in\mathbb{X}\\ C^{*}=-\mathbb{E}_{x}[D^{*}(x)]\end{cases}$
The sigmoid function $\sigma(z)$ is Lipschitz continuous(Rudin 2021) with constant $L_{\sigma}=\frac{1}{4}$. For any $D(x),D^{\prime}(x)$ and $C,C^{\prime}$, we have:
Then,
Taking the maximum over all $x$, we obtain
For the update of $C$, we have
Therefore, the joint mapping satisfies.
That is, the contraction constant is $\frac{1}{2}<1$.
Since the joint mapping is a contraction in the $(N+1)$-dimensional space(Banach 1922), by the Banach fixed-point theorem, the sequence $(\mathbf{D}_{n},C_{n})$ converges to a unique fixed point $(\mathbf{D}^{*},C^{*})$, regardless of the initial values.
Appendix C Training Details
@@
Table 4 presents the key configuration used for our Qwen2.5-7B experiment. Training was conducted on a single node with 8 A100 GPUs, and the model was trained for 110 steps using the veRL library (Sheng et al. 2024). For Qwen2.5-14B experiment, training was conducted on four nodes, each equipped with 8 A100 GPUs. The model was trained for 200 steps with a batch size of 256. Other configurations were kept consistent with those used in the Qwen2.5-7B experiment.
@@
Appendix D Evaluation Benchmarks
@@
  • MATH500 (Hendrycks et al. 2021b): A curated subset of the larger MATH dataset, consisting of 500 problems that span various mathematical fields such as algebra, geometry, and number theory.
@@
  • MinervaMath (Lewkowycz et al. 2022): A benchmark used to evaluate the performance of large language models on detailed, multi-step quantitative reasoning problems.
@@
  • OlympiadBench (He et al. 2024): A benchmark specifically designed for assessing models on Olympiad-level mathematics, including problems from national and international math competitions.
@@
  • GSM8K (Cobbe et al. 2021): A dataset of grade school math word problems, commonly used to test arithmetic and step-by-step reasoning in models.
Appendix E Effectiveness of Symmetric Sampling
As seen in Figure 7, after removing symmetric sampling, CDAS consistently outperforms the Random Sampling baseline. However, there is still an obvious decline in accuracy, especially in the later stages of training (after $100$ steps). We attribute this phenomenon to the accumulated estimation errors in model competence due to the imbalanced sampling. In the early stages of training, imbalanced sampling can cause the model’s competence to be either overestimated or underestimated. This discrepancy from the true competence accumulates as the number of training steps increases and hinders the convergence of CDAS, impairing the performance in the later stages.
Figure 7: Ablation study of symmetric sampling.
Appendix F Effectiveness of Warmup Phase
Our warm-up phase is specifically designed to address two potential criticisms. First, it directly mitigates the risk of inheriting and amplifying dataset biases. Without this initial calibration, the first few random batches could, by chance, overrepresent certain problem types (e.g., algebra in the MATH dataset). CDAS might then prematurely assign low difficulty to these common problems and high difficulty to unseen, underrepresented topics, effectively marginalizing them from future sampling. Our strategy prevents this by guaranteeing a comprehensive, initial assessment across the entire problem distribution, ensuring that a problem’s initial difficulty is based on an actual model attempt, not on an absence of data. Second, this approach enhances the algorithm’s generalizability and reduces its dependence on a dataset’s specific difficulty distribution. The warm-up acts as a mandatory “discovery phase,” allowing CDAS to construct a complete and fair map of the initial difficulty landscape before it starts making targeted decisions. This bootstraps the system with a high-quality, holistic foundation, ensuring that the subsequent efficiency gains from competence-difficulty alignment are built upon a stable and equitable understanding of the task.
The importance of this well-founded initial state is not merely practical; it is also theoretically motivated. Since CDAS is built on a fixed-point system, the choice of initial values can affect its convergence speed and stability. To investigate the impact of this deliberate initialization, we conduct an ablation study by removing the warm-up phase and instead starting with naive zero-initialized difficulties.
Figure 8: Ablation study of the warm-up phase
As observed in Figure 8, the training curve exhibits significant fluctuations in the early stages of training (the first $20$ steps) without the warmup stage. Specifically, in the first $6$ steps, the removal of the warmup phase leads to a notable decrease in response length along with a rapid increase in reward, suggesting that the model has over-learned from simpler samples. However, as the convergence of the entire fixed-point system is guaranteed, the two training curves gradually overlap, further corroborating the stability of our framework.